If the wage rate increases from $12 to $16 and, as a result, the quantity demanded of labor decreases from 20,000 workers to 16,000 workers, then the absolute value of the elasticity of demand for labor is ___________, which tells us that the demand for labor in this situation is ______________.

1.29; inelastic

0.78; elastic

0.78; inelastic

1.29; elastic
Answer #1

Answer
Elasticity of demand for labor=(change in quantity of labor/average quantityof labor)/(change in wage /average wage)
Change in quantity of labor=16000-20000=-4000
average quantity of labor=(16000+20000)/2=18000
change in wage=16-12=4
average wage=(16+12)/2=14
Elasticity of demand=(-4000/18000)/(4/14)
=-0.777777778
=-0.78
=0.78 ( absolute value)

the demand is inelastic as the elastcity is below 1.

Option 3